OpenStudy (anonymous):

Solve the following system of equations y = −x2 + 4 y = 2x + 1 Can someone help me with this one? I got no answer but I think it's wrong

Vocaloid (vocaloid):

-x^2 + 4 = 2x + 1 -x^2 - 2x + 3 = 0 factor

OpenStudy (anonymous):

would it be x = 1 or x = -3 ?

OpenStudy (abhisar):

Do you mean there are two equations \(\sf y = −x2 + 4~and~y = 2x + 1\) ?

OpenStudy (anonymous):

yeah

OpenStudy (abhisar):

You can subtract second equation from first one to get a quadratic equation in terms of x only and then solve it to find the value of x. Finally substitute this value of x in any of the parent equation to find the value of y

OpenStudy (anonymous):

So -x^2 +4 = 2x + 1 and I still get x=1 or -3

OpenStudy (abhisar):



OpenStudy (abhisar):

So that gives you a quadratic equation \(\sf x=\Large \frac{2 \pm \sqrt{2^2-(4 \times-1 \times 3)}}{2}\)

OpenStudy (abhisar):

Solve for c: